OpenStudy (bj11684):

Kylie opened a bank account. She deposited $76.50 into her account every month for 10 months. She used $30.50 every month to pay for singing lessons. After 10 months, she used 1 over 4 of the total money left in her account to go to a talent camp. What is the total amount of money Kylie spent to go to the talent camp? A.$97.00 B.$107.00 C.$115.00 D.$468.00 

TheSmartOne (thesmartone):

how much money did she deposit? 76.50 * 10 = ? how much money did she take out? 30.50 * 10 = ?

OpenStudy (bj11684):

76.50 * 10 = 3825

OpenStudy (bj11684):

30.50 * 10 = 305

OpenStudy (phi):

this does not look correct 76.50 * 10 = 3825

OpenStudy (robtobey2):

C.\[\frac{10}{4} (76.5\, -30.5)=115. \]

TheSmartOne (thesmartone):

@bj11684 try again: 76.50 * 10 = ? now, this number we get is how much she has in her bank account. We then need to subtract 305 from it for the Piano lessons. Now we multiply by 1/4 or just divide by 4 to get your final answer :)